First Generation College Attendance

Men

Women

Firsts

817.14

1118.86

Nonfirsts

1132.86

1551.14

Calculate Chi-squared and degree of freedom. Is there a significant difference at the p<0.05 level? Explain.
Answer #1

SOLUTION:

From given data,

First Generation College Attendance Men Women
Firsts 817.14 1118.86
Nonfirsts 1132.86 1551.14

Add up rows and columns:

First Generation College Attendance(O) Men Women Total
Firsts 817.14 1118.86 1936
Nonfirsts 1132.86 1551.14 2684
Total 1950 2670 4620

Calculate "Expected Value" for each entry:

Multiply each row total by each column total and divide by the overall total:

E = (Row total)(Column total) / Grand Total

E Men Women
Firsts 1936*1950 / 4620 =817.142857 1936*2670/4620 = 1118.857143
Nonfirsts 2684*1950/4620 = 1132.857143 2684*2670/4620 = 1551.1428571

Subtract expected from actual, square it, then divide by expected:

\chi ^{2} = ΣΙΟ-ΕΕ

Men Women
Firsts (817.14-817.142857)2 / 817.142857=0.0000000099 (1118.86-1118.857143)2/1118.857143 =0.0000000073
Nonfirsts (1132.86-1132.857143)2/1132.857143 = 0.0000000072 (1551.14-1551.1428571)2/1551.1428571 = 0.0000000053

Now add up those values:

\chi ^{2} =  ΣΙΟ-ΕΕ = 0.0000000099+ 0.0000000073+0.0000000072+0.0000000053 = 0.0000000297

Chi-Square = \chi ^{2} = 0.0000000297

Calculate Degrees of Freedom

= (rows − 1) \times (columns − 1)

DF = (2 − 1)(2 − 1) = 1×1 = 1

Critical value :

Critical value at  \alpha = 0.05 with   DF = 1

Хо.05- = 3.841

P- value

P- value at  \chi ^{2} = 0.0000000297 with   DF = 1

P- value = 0.999

Since test statistic is less than critical value, we fail to reject null hypothesis. At 5% level , i.e., there is no significant difference at the p<0.05 level.

Please thumbs-up / vote up this answer if it was helpful. In case of any problem, please comment below. I will surely help. Down-votes are permanent and not notified to us, so we can't help in that case.